Consider the 3–dimensional vector field F defined by F (x, y,z) = 12x2y2 +2z2 +1, 8x3y −3z, 4xz −3y −3.

(a) Write down the Jacobian matrix JF (x, y,z).

(b) Determine the divergence div F (x, y,z). (c) Determine curl F (x, y,z). (d) Give reasons why F has a potential function.

(e) Find a potential function of F.



1
Expert's answer
2021-06-01T17:36:46-0400

Ans:-

"F(x,y,z)=(12x ^2 y^ 2+2z^ 2+1,8x^ 3 y\u22123z,4xz\u22123y\u22123)"

or "\overrightarrow F=" "(12x ^2 y^ 2+2z^ 2+1)\hat i+(8x^ 3 y−3z)\hat j+(4xz−3y−3)\hat k"


"(i)" Jacobian Matrix "J(F(x,y,z)=\begin{bmatrix}
 \dfrac{df_x}{dx} & \dfrac{df_x}{dy}& \dfrac{df_x}{dz} \\\\
 \dfrac{df_y}{dx}& \dfrac{df_y}{dy}& \dfrac{df_y}{dz}\\\\
 \dfrac{df_z}{dx} & \dfrac{df_z}{dy}& \dfrac{df_z}{dz}
\end{bmatrix}"




="\begin{bmatrix}
 24xy^2& 24x^2y& 4z \\\\
 24xy^2 & 8x^3& -3\\\\
4z& -3&0
\end{bmatrix}"



"(ii)" "divF(x,y,z)"


"= (\dfrac{d}{dx}\hat{i}+\dfrac{d}{dy}\hat{j}+\dfrac{d}{dz}\hat{k}).((12x^2y^2+2z^2+1)\hat{i}+(8x^3y-3z)\hat{j}+(4xz-3y-3)\hat{k})"


"=24xy^2+8x^3+4x"


"(iii)" "curl F(x,y,z)=\begin{bmatrix}
 \hat{i} & \hat{j} &\hat{k}\\\\
 \dfrac{d}{dx} &\dfrac{d}{dy}&\dfrac{d}{dz} \\\\
 12x^2y^2+2z^2+1& 8x^3y-3z& 4xz-3y-3
\end{bmatrix}"






"=\hat i(−3−(−3))− \hat j(4z−4z)+\hat k(24x^ 2y−24x^ 2y)\\
=0"


"(iv)"

F is a potential function because, Whenever There exist a field, There exist a potential associated with it. We can calculate the potential function F by the integrating the given vector field.


"(v)" As we know -

Potential function is given by-

"F _v =\u2212\u222bF.dr"

"=\int ((12x^2y^2+2z^2+1)\hat{i}+(8x^3y-3z)\hat{j}+(4xz-3y-3)\hat{k})(dx\hat{i}+dy\hat{j}+dz\hat{z})"


"=(4x ^3y+2z^ 2x+x+4x^ 3y^ 2 \u22123yz+2x^ 2z\u22123xy\u22123x)"


"=8x ^3y^ 2+2x^ 2z+2xz^ 2\u22123xy\u22123yz\u22122x"